A limited has an authorized share capital of 200,000 shares of Kshs. 1 each out of which only 150,000 shares have been issued. Although the firm requested the shareholders to pay 80 cents per share, they were able to pay 50 cents per share.
Required:
Determine the:
i. Authorized share capital
ii. Issued share capital
iii. Called up share capital
iv. Uncalled up share capital
v. Paid up share capital
1
Expert's answer
2015-06-25T00:00:41-0400
A limited has an authorized share capital of 200,000 shares of Kshs. 1 each out of which only 150,000 shares have been issued. Although the firm requested the shareholders to pay 80 cents per share, they were able to pay 50 cents per share.
i. Authorized share capital is 200,000*1 = 200,000 Kshs.
ii. Issued share capital is 150,000*1 = 150,000 Kshs.
iii. Called up share capital is 150,000*0.8 = 120,000 Kshs.
iv. Uncalled up share capital is 150,000*(1 - 0.8) = 30,000 Kshs.
v. Paid up share capital is 150,000*0.5 = 75,000 Kshs.